7.What is the common difference in the sequence3,5,7,9,11,⋯
a.-2
b.2
c.3
d.-3

⎛1 2 3⎞
8.Calculate the determinant of the matrix A= ⎜ 4 5 0⎟
⎝2 1 4⎠
a.

⎛ 20 −16 −6 ⎞
A= ⎜−5 −2 3 ⎟
⎝−13 12 −41 ⎠

b.
⎛ 20 −16 −6 ⎞
A = ⎜−5 −2 3 ⎟
⎝−13 12 −3 ⎠

c.
⎛ 20 −16 −6 ⎞
A = ⎜−5 −2 3 ⎟
⎝−15 12 −3 ⎠

d.
⎛ 20 −12 −6 ⎞
A = ⎜−5 −2 3 ⎟
⎝−13 12 −41 ⎠
1
Expert's answer
2019-04-02T11:37:13-0400

7). d=5-3=2.

Answer b.

8). |A|=1*5*4+3*4*1+0*2*2-3*5*2-2*4*4-1*1*0=-30.